OpenStudy (anonymous):

-3x+y=7 3x+2y=2 Eliminate. Show me how please. :D

OpenStudy (anonymous):

you can simply add them and get 3y=9, then solve for y

OpenStudy (anonymous):

But then how do you get what x =?

OpenStudy (anonymous):

Submit back the y to find x=-4/3

OpenStudy (anonymous):

y = 9/3 = 3 Hence replace y in one of your original equations and you'll get x Let's do it on first equation. -3x + y = 7 -3x = 7 -y x = -( 7 -y)/3 x = -( 7 - 3)/3 x = -4/3

OpenStudy (anonymous):

Thank you